# When higher-order interactions enhance synchronization: the case of the Kuramoto model

### v1

@@ Methods
In this analysis, we fix $N_{0}=10$ nodes, $N_{1}=20$ links, and $N_{2}=10$ triangles for the hypergraphs to be generated. Then, we define the coupling strength ranges for pairwise and higher-order interactions, $K_{1}$ and $K_{2}$, both spanning from 0 to $0.6$. A grid of $51\times 51$ points is created over these ranges, and for each grid point, we integrate (1) (or (3), with results shown in the Supplementary Materials [^9] For this case, the range of the coupling strength $K_{2}$ spans from 0 to $1$, as the values of $K_{2}$ maximizing the averaged order parameter $R$ are slightly higher than for the case shown in the Main Text.) over $N_{\mathcal{H}}=300$ randomly generated hypergraph with the prescribed $(N_{0},N_{1},N_{2})$, ensuring either either $\mathcal{H}$-connectivity or $1$-connectivity. For each hypergraph, frequencies $\omega_{j}$ and initial phases $\vartheta_{0,j}$, for $j=1,\dots,N_{0}$, are drawn uniformly from $[0,0.3]$ and $0,\vartheta_{0}^{\mathrm{max}}]$, respectively, with $\vartheta_{0}^{\mathrm{max}}\in\{0.1,2\pi\}$. As noted earlier, the choice of frequency distribution does not qualitatively affect the results. The large number of random hypergraphs, frequency sets, and initial conditions mitigates finite-size effects. Increasing $N_{\mathcal{H}}$ further was tested and showed no significant impact. For each integration, we compute the Kuramoto order parameter $R$ and the Kuramoto-Daido order parameter $R_{2}$, averaging their values over the the hypergraph samples at each grid point as described above. When starting from incoherent initial conditions (i.e., $\vartheta_{0}^{\mathrm{max}}=2\pi$), we also mark, for each $K_{1}$, the value of $K_{2}$ yielding the maximum order parameters.
In Figure 3, we show the order parameter $R$ obtained for the higher-order Kuramoto in (1) starting from both coherent ($\vartheta_{0}^{\mathrm{max}}=0.1$) and incoherent ($\vartheta_{0}^{\mathrm{max}}=2\pi$) initial conditions. We also performed the analysis with different numbers of links and triangles, namely $N_{1}=40$, $N_{2}=6$ and $N_{1}=15$, $N_{2}=6$, observing no meaningful difference (results not shown).
In the Supplementary Materials, we show analogous results for the Kuramoto-Daido order parameter $R_{2}$ of the model in (1), as well as $R$ and $R_{2}$ for the model in (3). All configurations exhibit the same qualitative behavior shown in the Main Text.
@@
The experimental conditions were generated by varying coupling strengths $K_{1},K_{2}$ in (1) and the allocation preference $b_{1}$. Our goal was to determine the optimal hyperedge allocation for a given physical regime, as characterized by the coupling strengths. To this end, we varied their value while keeping their sum fixed: we selected $K_{1}\in[0,K^{\mathrm{sum}}]$, with $K^{\mathrm{sum}}=0.3$, using $51$ linearly spaced points, and consequently set $K_{2}=K^{\mathrm{sum}}-K_{1}$. The allocation parameter $b_{1}$ was varied over $[0,1]$, also using $51$ linearly spaced points, resulting in $51\times 51=2601$ experimental conditions. For each value of $b_{1}$, we generated $N_{\mathcal{H}}=500$ random hypergraphs and sampled random initial conditions $\vartheta_{0}$ uniformly from $0,\vartheta_{0}^{\mathrm{max}}]^{N_{0}}$, as well as natural frequencies uniformly from $0,0.3]^{N_{0}}$, with $\vartheta_{0}^{\mathrm{max}}$ selected as either $\pi/2$ or $2\pi$. When generating the hypergraphs, the $N_{1}$ and $N_{2}$ hyperedges are selected randomly from all possible ones, each with equal probability. If the resulting hypergraph is not $\mathcal{H}$-connected, it is discarded and a new one is generated in its place, repeating this process until an $\mathcal{H}$-connected hypergraph is obtained.
At this point, a grid of $2601$ experimental conditions are obtained, each with $500$ realizations of hypergraph structure, initial conditions, and natural frequencies. For each realization, we simulate the Kuramoto model (4) and record the average order parameter $R$ in the second half of each simulation. The value of $R$ obtained over all $500$ realizations is averaged to obtain a single value for each experimental condition, thus creating one of the panels in Figure 4. The different panels were created by changing the value of the cost associated with $2$-hyperedges, $c_{2}$, and the range of the initial phase conditions, $\vartheta_{0}^{\mathrm{max}}$, to produce more or less homogeneous initial states. In total $2601\times 500\times 6=7.803.000$ were run to realize Figure 4.
@@ Acknowledgments
The work of R.M. is supported by a JSPS postdoctoral fellowship, grant 24KF0211. H.N. acknowledges JSPS KAKENHI 25H01468, 25K03081, and 22H00516 for financial support. The authors are grateful to Daniele Proverbio for discussions and feedback on the work. R.M. and M.C. would like to thank Claudio Altafini and all the organizers of the ELLIIT Focus Period on Network Dynamics and Control at Linköping University in September 2023, where this collaboration started.
@@
Supplementary Material
In the Results section of the Main Text, we presented numerical analyses of the higher-order Kuramoto model with higher-order interaction of the form $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$ (for the dynamics of the $j$-th oscillator). In this Supplementary file, we additionally show results for the model featuring higher-order interactions of the form $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$, with a vector form model derived in the next Section. We also provide, for both types of higher-order interaction, the value of the Kuramoto-Daido order parameter $R_{2}$ obtained in the simulations.
@@
Figure Discussed in Higher-order interaction Metric
Fig. 3 Main Text $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$ $R$
Fig. SM1 Sec. Results with a different higher-order interaction $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$ $R$
Fig. SM3 Sec. Results for the Kuramoto-Daido second order parameter $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$ $R_{2}$
Fig. SM4 Sec. Results for the Kuramoto-Daido second order parameter $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$ $R_{2}$
@@
Figure Discussed in Higher-order interaction Metric $\omega^{\mathrm{max}}$
Fig. 4 Main Text $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$ $R$ 0.3
Fig. SM2 Sec. Results with a different higher-order interaction $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$ $R$ 0.3
Fig. SM5 Sec. Results for the Kuramoto-Daido second order parameter $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$ $R_{2}$ 0.3
Fig. SM6 Sec. Results for the Kuramoto-Daido second order parameter $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$ $R_{2}$ 0.3
@@
In the Main Text, we analyzed the higher-order Kuramoto model with interactions of the form $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$, which naturally arise from phase reduction León et al. (2024). However, other authors have considered a different harmonic for the higher-order interactions, namely $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$ Skardal and Arenas (2020), which was recently shown to also emerge from phase reduction León et al. (2025). We complemented the analysis in the Main Text by performing analogous simulations for the Kuramoto model with this alternative higher-order interaction term. To enable more efficient simulations, we reformulated the model in vector form, as detailed below. The results, shown in Figures SM1, SM2, are qualitatively consistent to those shown in the Main Text, with the following differences. In the coupling strength analysis (Figure SM1), the higher-order interaction coupling strength ($K_{2}$) that maximizes synchronization is slightly larger. In the hyperedge allocation analysis, triangles are less effective at promoting synchronization, which slightly modifies the shape of the black line (indicating optimal allocations) compared to Figure 4 in the Main Text.
Vector form of the higher-order Kuramoto model with term $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$
Here, we derive a vector form for the Kuramoto model with higher-order interactions given by $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$. As in the Main Text, the model can be written in vector form as
$\dot{\vartheta}=\omega+\frac{K_{1}}{\langle d^{(1)}\rangle}B_{1}\sin\left(-(B_{1})^{\mathsf{T}}\vartheta\right)+\frac{K_{2}}{2\langle d^{(2)}\rangle}Y^{(2)}_{\mathrm{s}}\sin\left(-(Y^{(2)}_{\mathrm{a}})^{\mathsf{T}}\vartheta\right),$ (SM1)
but with different matrices $Y^{(2)}_{\mathrm{a}}$ and $Y^{(2)}_{\mathrm{s}}$. Namely, let $z^{(2)}\in\mathbb{R}^{N_{0}}$ be the vector collecting all coupling terms of the form $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$ Skardal and Arenas (2020) associated with $2$-hyperedges, with the $j$-th component given by
$z_{j}^{(2)}\coloneqq\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j}).$
@@
Then, it holds that $z^{(2)}=Y^{(2)}_{\mathrm{s}}\sin(-(Y^{(2)}_{\mathrm{b}})^{\mathsf{T}}\vartheta)$, which confirms the form in Eq. (SM1).
@@
In Figures SM3, SM4, SM5, and SM6, we show the value of the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\vartheta_{j}}\right\rvert$ obtained for the simulations considered, with both kinds of higher-order interactions. Comparing these results with those in the Main Text and the previous Section, we observe that $R_{2}$ is large in the same regions where $R$ is large, indicating clustering phenomena are absent, on average.
@@
Figure SM1: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
@@
Figure SM2: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\vartheta_{k}-\vartheta_{l}-\vartheta_{j})$. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
@@
Figure SM3: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\vartheta_{k}+\vartheta_{l}-2\vartheta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$ (see Methods). Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.

### v2

@@ Methods
In this analysis, we fix $N_{0}=10$ nodes, $N_{1}=20$ links, and $N_{2}=10$ triangles for the hypergraphs to be generated. Then, we define the coupling strength ranges for pairwise and higher-order interactions, $K_{1}$ and $K_{2}$, both spanning from 0 to $0.6$. A grid of $51\times 51$ points is created over these ranges, and for each grid point, we integrate (1) (or (2), with results shown in the Supplementary Materials [^10] For this case, the range of the coupling strength $K_{2}$ spans from 0 to $1$, as the values of $K_{2}$ maximizing the averaged order parameter $R$ are slightly higher than for the case shown in the Main Text.) over $N_{\mathcal{H}}=300$ randomly generated hypergraph with the prescribed $(N_{0},N_{1},N_{2})$, ensuring either either $\mathcal{H}$-connectivity or $1$-connectivity. For each hypergraph, frequencies $\omega_{j}$ and initial phases $\theta_{0,j}$, for $j=1,\dots,N_{0}$, are drawn uniformly from $[0,0.3]$ and $0,\theta_{0}^{\mathrm{max}}]$, respectively, with $\theta_{0}^{\mathrm{max}}\in\{0.1,2\pi\}$. As noted earlier, the choice of frequency distribution does not qualitatively affect the results. The large number of random hypergraphs, frequency sets, and initial conditions mitigates finite-size effects. Increasing $N_{\mathcal{H}}$ further was tested and showed no significant impact. For each integration, we compute the Kuramoto order parameter $R$ and the Kuramoto-Daido order parameter $R_{2}$, averaging their values over the the hypergraph samples at each grid point as described above. When starting from incoherent initial conditions (i.e., $\theta_{0}^{\mathrm{max}}=2\pi$), we also mark, for each $K_{1}$, the value of $K_{2}$ yielding the maximum order parameters.
In Figure 3, we show the order parameter $R$ obtained for the higher-order Kuramoto in (1) starting from both coherent ($\theta_{0}^{\mathrm{max}}=0.1$) and incoherent ($\theta_{0}^{\mathrm{max}}=2\pi$) initial conditions. We also performed the analysis with different numbers of links and triangles, namely $N_{1}=40$, $N_{2}=6$ and $N_{1}=15$, $N_{2}=6$, observing no meaningful difference (results not shown).
In the Supplementary Materials, we show analogous results for the Kuramoto-Daido order parameter $R_{2}$ of the model in (1), for normally distributed frequencies, for larger structures, as well as $R$ and $R_{2}$ for the model in (2). All configurations exhibit the same qualitative behavior shown in the Main Text.
@@
The experimental conditions were generated by varying coupling strengths $K_{1},K_{2}$ in (1) and the link allocation fraction $b_{1}$. Our goal was to determine the optimal hyperedge allocation for a given physical regime, as characterized by the coupling strengths. To this end, we varied their value while keeping their sum fixed: we selected $K_{1}\in[0,K^{\mathrm{sum}}]$, with $K^{\mathrm{sum}}=0.3$, using $51$ linearly spaced points, and consequently set $K_{2}=K^{\mathrm{sum}}-K_{1}$. The allocation parameter $b_{1}$ was varied over $[0,1]$, also using $51$ linearly spaced points, resulting in $51\times 51=2601$ experimental conditions. For each value of $b_{1}$, we generated $N_{\mathcal{H}}=500$ random hypergraphs and sampled random initial conditions $\theta_{0}$ uniformly from $0,\theta_{0}^{\mathrm{max}}]^{N_{0}}$, as well as natural frequencies uniformly from $0,0.3]^{N_{0}}$, with $\theta_{0}^{\mathrm{max}}$ selected as either $\pi/2$ or $2\pi$. When generating the hypergraphs, the $N_{1}$ and $N_{2}$ hyperedges are selected randomly from all possible ones, each with equal probability. If the resulting hypergraph is not $\mathcal{H}$-connected, it is discarded and a new one is generated in its place, repeating this process until an $\mathcal{H}$-connected hypergraph is obtained.
At this point, a grid of $2601$ experimental conditions are obtained, each with $500$ realizations of hypergraph structure, initial conditions, and natural frequencies. For each realization, we simulate the Kuramoto model (3) and record the average order parameter $R$ in the second half of each simulation. The value of $R$ obtained over all $500$ realizations is averaged to obtain a single value for each experimental condition, thus creating one of the panels in Figure 5. The different panels were created by changing the value of the cost associated with $2$-hyperedges, $c_{2}$, and the range of the initial phase conditions, $\theta_{0}^{\mathrm{max}}$, to produce more or less homogeneous initial states. In total $2601\times 500\times 6=7{,}803{,}000$ simulations were run to realize Figure 5. Also for this analysis, in the Supplementary Materials, we show analogous results for the Kuramoto-Daido order parameter $R_{2}$ of the model in (1), for normally distributed frequencies, for larger structures, as well as $R$ and $R_{2}$ for the model in (2). All configurations exhibit the same qualitative behavior shown here.
@@ Acknowledgments
R.M. and M.C. contributed equally to this work. R.M. acknowledges JSPS KAKENHI 24KF0211 for financial support. H.N. acknowledges JSPS KAKENHI 25H01468, 25K03081, and 22H00516 for financial support. The authors are grateful to Daniele Proverbio for discussions and feedback on the work. R.M. and M.C. would like to thank Claudio Altafini and all the organizers of the ELLIIT Focus Period on Network Dynamics and Control held at Linköping University in September 2023, where this collaboration started.
@@
Supplementary Materials
In the Results section of the Main Text, we presented numerical analyses of the higher-order Kuramoto model with higher-order interaction of the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$ (for the dynamics of the $j$-th oscillator), which we refer to as the (1,1,2)(1,1,-2) interaction; see Eq. (1) in the Main Text. In this Supplementary file, we additionally show results for the model featuring higher-order interactions of the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$—which we refer to as the (2,1,1)(2,-1,-1) interaction (see Eq. (2) in the Main Text)—with a vector form model derived in § Results with a different higher-order interaction. For both types of higher-order interaction, we also provide the value of the Kuramoto-Daido order parameter $R_{2}$ obtained in the simulations (§ Results for the Kuramoto-Daido second order parameter), as well as the simulation results obtained with frequencies distributed normally (§ Results for Gaussian frequency distribution). In § Results for larger hypergraphs, we provide results obtained for larger structures. All results confirm the findings of the Main Text.
@@
Figure Discussed in # Nodes Higher-order interaction Metric frequency distribution
Figs. 3, 4 Main Text 10 (1,1,2)(1,1,-2) $R$ uniform
Fig. SM1 Sec. Results with a different higher-order interaction 10 (2,1,1)(2,-1,-1) $R$ uniform
Fig. SM3 Sec. Results for the Kuramoto-Daido second order parameter 10 (1,1,2)(1,1,-2) $R_{2}$ uniform
Fig. SM4 Sec. Results for the Kuramoto-Daido second order parameter 10 (2,1,1)(2,-1,-1) $R_{2}$ uniform
Figs. SM7, SM8 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R$ Gaussian
Fig. SM9 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R$ Gaussian
Fig. SM10 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R_{2}$ Gaussian
Fig. SM11 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R_{2}$ Gaussian
Fig. SM16 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R$ uniform, Gaussian
Fig. SM17 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{2}$ uniform, Gaussian
@@
Figure Discussed in # Nodes Higher-order interaction Metric frequency distribution
Fig. 5 Main Text 10 (1,1,2)(1,1,-2) $R$ uniform
Fig. SM2 Sec. Results with a different higher-order interaction 10 (2,1,1)(2,-1,-1) $R$ uniform
Fig. SM5 Sec. Results for the Kuramoto-Daido second order parameter 10 (1,1,2)(1,1,-2) $R_{2}$ uniform
Fig. SM6 Sec. Results for the Kuramoto-Daido second order parameter 10 (2,1,1)(2,-1,-1) $R_{2}$ uniform
Fig. SM12 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R$ Gaussian
Fig. SM13 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R$ Gaussian
Fig. SM14 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R_{2}$ Gaussian
Fig. SM15 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R_{2}$ Gaussian
Fig. SM18 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R$ uniform
Fig. SM19 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{2}$ uniform
Fig. SM20 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{1}$ Gaussian
Fig. SM21 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{2}$ Gaussian
@@
In the Main Text, we analyzed the higher-order Kuramoto model with (1,1,2)(1,1,-2) higher-order interactions. However, from phase reduction, both second-order pairwise León and Pazó (2019) and first-order many-body León et al. (2025), also the (2,1,1)(2,-1,-1) higher-order interaction can emerge. For this reason,
we complemented the analysis in the Main Text by performing analogous simulations for the Kuramoto model with this alternative higher-order interaction term. To enable more efficient simulations, we reformulated the model in vector form, as detailed below in § Vector form of the higher-order Kuramoto model with the (2,1,1)(2,-1,-1) interaction. The results, shown in Figures SM1, SM2, are qualitatively consistent to those shown in the Main Text, with the following differences. In the coupling strength analysis (Figure SM1), the higher-order interaction coupling strength ($K_{2}$) that maximizes synchronization is slightly larger. In the hyperedge allocation analysis, triangles are less effective at promoting synchronization, which slightly modifies the shape of the black line (indicating optimal allocations) compared to Figure 4 in the Main Text.
Vector form of the higher-order Kuramoto model with the (2,1,1)(2,-1,-1) interaction
Here, we derive a vector form for the Kuramoto model with (2,1,1)(2,-1,-1) higher-order interactions. As in the Main Text, the model can be written in vector form as
$\dot{\theta}=\omega+\frac{K_{1}}{\langle d^{(1)}\rangle}B_{1}\sin\left(-(B_{1})^{\mathsf{T}}\theta\right)+\frac{K_{2}}{2\langle d^{(2)}\rangle}Y^{(2)}_{\mathrm{s}}\sin\left(-(Y^{(2)}_{\mathrm{a}})^{\mathsf{T}}\theta\right),$ (SM1)
but with different matrices $Y^{(2)}_{\mathrm{a}}$ and $Y^{(2)}_{\mathrm{s}}$. Namely, let $z^{(2)}\in\mathbb{R}^{N_{0}}$ be the vector collecting all coupling terms of the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$ Skardal and Arenas (2020) associated with $2$-hyperedges, with the $j$-th component given by
$z_{j}^{(2)}\coloneqq\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(2\theta_{k}-\theta_{l}-\theta_{j}).$
@@
Then, it holds that $z^{(2)}=Y^{(2)}_{\mathrm{s}}\sin(-(Y^{(2)}_{\mathrm{b}})^{\mathsf{T}}\theta)$, which confirms the form in Eq. (SM1).
@@
In Figures SM3, SM4, SM10, SM11, SM17 (for the coupling strength analysis) and SM5, SM6, SM14, SM15, SM19, SM21 (for the hyperedge allocation analysis), we show the value of the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\theta_{j}}\right\rvert$ obtained for the simulations considered, with both kinds of higher-order interactions. Comparing these results with those in the Main Text and in this Supplementary Materials, we observe that $R_{2}$ is large in the same regions where $R$ is large, indicating clustering phenomena are absent, on average.
Results for Gaussian frequency distribution
In Figures SM7, SM8, SM9, SM10, SM11, SM16, SM17 (for the coupling strength analysis) and SM12, SM13, SM14, SM15, SM20, SM21 (for the hyperedge allocation analysis), we show the order parameter $R$ and the Kuramoto-Daido order parameter $R_{2}$ obtained for all the simulations considered, in the case that, for each hypergraph realization, oscillator frequencies $\omega_{j}$ are drawn randomly from a Gaussian distribution with mean $0.15$ and standard deviation $0.1$.
The results are qualitatively analogous to those reported in the Main Text, where a uniform frequency distribution was considered instead.
Results for larger hypergraphs
In Figures SM16, SM17 (for the coupling strength analysis) and SM18, SM19, SM20, SM21 (for the hyperedge allocation analysis), we show the value of the order parameter $R$ and the Kuramoto-Daido order parameter $R_{2}$ obtained for simulations of larger hypergraphs of $N_{0}=100$, with both uniform and normal frequency distributions.
To illustrate the choice of the simulation parameters, we recall the expressions for (i) the minimum number $N_{1}^{\mathrm{min}}$ of links required for $1$-connectivity, (ii) the minimum number $N_{2}^{\mathrm{min}}$ of triangles required for $2$-connectivity , (iii) the maximum number $N_{1}^{\mathrm{max}}$ of distinct links, and (iv) the maximum number $N_{2}^{\mathrm{max}}$ of distinct triangles. For a structure of $N_{0}$ nodes, these are
$N_{1}^{\mathrm{min}}(N_{0})=N_{0}-1,\qquad N_{2}^{\mathrm{min}}(N_{0})=\left\lceil\frac{N_{0}-1}{2}\right\rceil,\qquad N_{1}^{\mathrm{max}}(N_{0})=\frac{N_{0}(N_{0}-1)}{2},\qquad N_{2}^{\mathrm{max}}(N_{0})=\frac{N_{0}(N_{0}-1)(N_{0}-2)}{6}.$
For the coupling strength analysis, we set the numbers of links $N_{1}$ and triangles $N_{2}$ to as to preserve the same interaction densities used for the smaller hypergraphs ($N_{0}=10$). Preserving link and triangle densities requires keeping the ratios $N_{1}/N_{1}^{\mathrm{max}}$ and $N_{2}/N_{2}^{\mathrm{max}}$ constant. For $N_{0}=10$, we have $N_{1}^{\mathrm{max}}(10)=45$ and $N_{2}^{\mathrm{max}}(10)=120$, and we set $N_{1}=20$ and $N_{2}=10$. Then, for $N_{0}=100$, we compute $N_{1}^{\mathrm{max}}(100)=4950$ and $N_{2}^{\mathrm{max}}(100)=161{,}700$; thus, matching the previous densities yields $N_{1}=2200$ and $N_{2}=13{,}475$.
Owing to the very high computational cost associated with such a large number of triangles, in Figs. SM16 and SM17 we adopted a coarser $4\times 7$ parameter grid and a reduced number of hypergraph realizations $N_{\mathcal{H}}=100$. In this case, the higher-order Kuramoto model was integrated for 1000 time units (instead of $600$), as convergence required longer times (convergence was further verified by extending the simulations up to $3000$ time units, with no qualitative changes observed). Despite the coarser grid, the results confirm the trends observed for smaller structures.
For choosing the budget value $J$ for the hyperedge allocation analysis, we proceeded as follows. Let $c_{1}^{\mathrm{min}}$ (resp, $c_{2}^{\mathrm{min}}$) be the minimum cost associated to links (resp. triangles), and let $c_{1}^{\mathrm{max}}$ (resp, $c_{2}^{\mathrm{max}}$) denote the corresponding maximum cost. In our study, $c_{1}^{\mathrm{min}}=c_{1}^{\mathrm{max}}=1$, $c_{2}^{\mathrm{min}}=1$, and $c_{2}^{\mathrm{max}}=5$. The budget $J$ was then chosen to satisfy two conditions: (i) $J\geq\max\{c_{1}^{\mathrm{max}}N_{1}^{\mathrm{min}},c_{2}^{\mathrm{max}}N_{2}^{\mathrm{min}}\}$, ensuring $\mathcal{H}$-connected hypergraphs could always be generated; (ii) $J\leq\min\{c_{1}^{\mathrm{min}}N_{1}^{\mathrm{max}},c_{2}^{\mathrm{min}}N_{2}^{\mathrm{max}}\}$, so that no portion of the budget is wasted (see Methods for details). For $N_{0}=100$, we have $N_{1}^{\mathrm{min}}(100)=99$, $N_{2}^{\mathrm{min}}(100)=50$, $N_{1}^{\mathrm{max}}(100)=4950$, and $N_{2}^{\mathrm{max}}(100)=161{,}700$. Hence, conditions (i) and (ii) yield $J\geq\max\{1\cdot 99,5\cdot 50\}=250$ and $J\leq\max\{1\cdot 4950,1\cdot 161{,}700\}=4950$. This defines the admissible range $J\in[250,4950]$, from which we selected $J=600$. In this case, we used a $31\times 31$ parameter grid and $N_{\mathcal{H}}=100$ hypergraph realizations. The qualitative findings reported in the Main Text are confirmed also in this setting.
@@
Figure SM1: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
@@
Figure SM2: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
@@
Figure SM3: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$ (see Methods). Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.